# SETA: Semantic-Aware Token Augmentation for Domain Generalization

### v1

@@ Abstract
Domain generalization (DG) aims to enhance the model robustness against domain shifts without accessing target domains. A prevalent category of methods for DG is data augmentation, which focuses on generating virtual samples to simulate domain shifts. However, existing augmentation techniques in DG are mainly tailored for convolutional neural networks (CNNs), with limited exploration in token-based architectures, i.e., vision transformer (ViT) and multi-layer perceptrons (MLP) models. In this paper, we study the impact of prior CNN-based augmentation methods on token-based models, revealing their performance is suboptimal due to the lack of incentivizing the model to learn holistic shape information. To tackle the issue, we propose the SEmantic-aware Token Augmentation (SETA) method. SETA transforms token features by perturbing local edge cues while preserving global shape features, thereby enhancing the model learning of shape information. To further enhance the generalization ability of the model, we introduce two stylized variants of our method combined with two state-of-the-art style augmentation methods in DG. We provide a theoretical insight into our method, demonstrating its effectiveness in reducing the generalization risk bound. Comprehensive experiments on five benchmarks prove that our method achieves SOTA performances across various ViT and MLP architectures. Our code is available at https://github.com/lingeringlight/SETA.
@@ I Introduction
Among recent DG methods, data augmentation methods have exhibited promising performance [19, 11, 20, 13, 14]. These methods primarily assume that domain gaps are caused by the disparity in style, thus introducing channel-level style perturbations for simulating potential domain shifts. While these methods have achieved notable success on CNN architectures, limited research investigates their effects on ViT or MLP models. Existing ViT and MLP models have shown excellent generalization ability [13, 21, 22]. Some pioneering works have pointed out that it is attributed to their ability to learn spatial dependencies among tokens, which encode structural shape information of objects, thus exhibiting a superior shape bias than CNNs [15, 13]. However, existing DG augmentation methods mainly perturb style information at the channel level, lacking explicit enhancement of shape information at the spatial level. Moreover, some studies have revealed the presence of spurious edges [23], e.g., category-irrelevant local edges. Spurious edges are not encoded in style information, yet severely affect the model learning of domain-invariant shape features [24]. Therefore, whether existing style augmentation DG methods could enhance the shape bias of ViT or MLP models remains an open question.
@@
In this paper, we initially analyze the impact of existing SOTA augmentation methods in DG on the shape bias of the models. Specifically, we evaluate the robustness of these methods to diverse biased perturbations, focusing on two representative advanced methods: DSU [14], which is based on statistical perturbation, and ALOFT [13], which exploits frequency space augmentation. The experiments are conducted on the representative MLP-like model GFNet [18]. As shown in Fig. 1 (a), both DSU and ALOFT methods exhibit the ability to enhance the model generalization. The enhancement is attributed to their perturbations on style features, implicitly boosting the ability of the model to capture edge information, as indicated in Fig. 1 (b). Moreover, the enhanced edge-capturing ability also contributes to the increased model robustness against style perturbations, as depicted in Fig. 1 (c). However, it is crucial to note that edge information encompasses not only global object shapes but also local spurious edges that are domain-specific [23]. Since existing style augmentation methods for DG lack spatial token perturbation, they fail to suppress the domain-specific local edges, resulting in the limited ability of the model to perceive the holistic shape structures of objects, as proved in Fig. 1 (d). Consequently, the model inevitably overfits the tokens encoding domain-specific features, which severely impedes its generalizability to unseen domains. These observations prompt a critical question: Could we augment edge features at the token level to explicitly enhance shape bias of the model, thus further improving the generalization ability of ViT and MLP architectures.
@@
To this end, we propose a novel SEmantic-aware Token Augmentation (SETA) for DG, aimed at enhancing the generalization performance of ViT and MLP models by introducing token-level local edge perturbations. Unlike previous DG augmentation methods that focus on diversifying channel-level styles, our method concentrates on generating token-level edge perturbations to guide the model in emphasizing global shape features during training. Specifically, we initially design the Energy-based Edge Tokens Selection (ETS) module, which extracts edge maps from input features using low-pass filters and calculates the energy associated with each token to distinguish the edge-relevant tokens. Subsequently, random pairs of samples are selected from the batch, with one sample retaining only the edge tokens and the other being modified by the Shape Tokens Shuffling (STS) module that shuffles token locations within the sample. Finally, we conduct the random mixing of the two samples using either Mixup or CutMix, with the augmented sample being assigned the label of the sample providing edge tokens. Furthermore, we extend our method by integrating SOTA style augmentation techniques to enhance both shape sensitivity and robustness to style shifts. We evaluate our methods on five challenging datasets, employing various ViT and MLP architectures. The results prove the effectiveness of our method to enhance model generalization. Our contributions are summarized as follows:
@@
  • We theoretically prove that the incentivizing shape sensitivity through our SETA could tighten the generalization risk bound and improve the generalizability of the model.
@@ II Related Work
Domain generalization. Domain generalization (DG) aims to learn generalized representations from source domains that generalize well to arbitrary unseen target domains. Existing DG methods address domain shift from various perspectives, including domain alignment [25, 26], meta-learning [9, 27], data augmentation [28, 29], ensemble learning [30, 31] and regularization methods [32]. The majority of relevant works to our work address the DG issue by data augmentation. These methods operate under the assumption that domain gaps predominantly arise from style shifts, thereby endeavoring to simulate domain shifts by generating diverse styles while keeping semantics consistency. Specifically, one prevalent line is to perturb the feature statistics (i.e., mean and variance) by sample interpolation [33, 19], modeling distribution [34, 14] or learning methods [35, 36]. Moreover, some recent works propose to leverage the semantic-preserving property of the Fourier phase component and introduce diverse noise into the amplitude spectrum to generate stylized data [37, 38]. However, these methods are mainly designed for CNN architectures, which focus on perturbing channel-level style features. When applied to ViT or MLP models that learn spatial dependencies among different tokens, as these methods lack token-level augmentation, the model inevitably overfits domain-specific tokens that encode spurious local edge information, which impedes its generalizability. Therefore, we propose a token-level augmentation method to adaptively extract edge tokens and introduce local texture noise, thereby enhancing the model’s ability to capture shape information and boosting its generalization to unseen target domains.
@@
Figure 2: An overview of the proposed SETA. Our SETA consists of three core modules, including 1) Energy-based Edge Tokens Selection (ETS) that distinguishes and extracts tokens that contain edge information; 2) Shape Tokens Shuffling (STS) that generates texture noise by shuffling tokens from another sample, which disrupts holistic shape while keeping local edges; 3) Token Mixing module that superposes edge tokens from an object sample onto the shuffled tokens from another sample. The augmented sample is assigned the label of the sample contributing the edge tokens. We design two stylized variants of SETA, utilizing SOTA DG augmentation methods, i.e., DSU [14] and ALOFT [13], to create stylized token-shuffled samples to simulate potential domain shifts.
@@
Data augmentation for improving shape bias. Previous studies have revealed that compared to human vision, CNNs are revealed to be more sensitive to image texture for object classification, indicating that the trained models lack a full understanding of different object categories [39, 40]. To reduce this disparity, augmentation methods aim to perturb texture representations, which could explicitly reduce the texture bias and implicitly boost the shape bias of the model [40]. Several works also indicate that an increased shape bias in CNNs is associated with improved model robustness [41, 42]. The exploration of texture and shape biases has extended to ViTs and MLPs, showing that they exhibit a lower texture bias compared to CNNs [43, 15]. In DG scenarios, the texture bias of the model has greatly damaged its ability to generalize across different domains [44, 45]. To tackle the issue, many DG techniques diversify data by perturbing style information while keeping semantic elements (e.g., edges and shapes) consistent [19, 37]. However, these methods ignore that spurious edges (i.e., shape-irrelevant edges) may exist in the remaining features except style and texture features, which potentially limit their ability to enhance the shape bias of the model [23]. In the paper, we introduce a semantic-aware token augmentation method, which could discern and perturb spurious edges while retaining crucial shape information. Our method could compel the model to learn a comprehensive understanding of object shape from the edge-perturbed representations, thereby enhancing its shape bias effectively.
@@ III-A Setting and Overview
In standard DG model training, the model is directly trained to predict the label of the original samples. However, as emphasized in [46, 39, 42], this training method could induce sensitivity to style and texture in source domains, consequently hindering model generalization on unseen target domains. To tackle this issue, we propose a novel token-level augmentation method, Semantic-aware Token Augmentation (SETA), tailored for both ViT and MLP models. The key idea of SETA is to generate shape-augmented representations by mixing edge-related tokens within a sample and shape-disrupted tokens from other samples. This operation compels the model to distinguish the overall shape of objects from the perturbed representations, thereby enhancing its sensitivity to shape information. To further enhance the model robustness to domain shifts, we explore a paradigm that integrates SETA with SOTA augmentation methods in DG, yielding two stylized variants of SETA. An overview of our method is depicted in Fig. 2. Below we introduce the main components of our method and provide a theoretical insight into its effectiveness.
III-B Energy-based Edge Tokens Selection
To extract feature tokens encoding edge information, we design the Energy-based Edge Tokens Selection (ETS) module. The ETS module extracts edge maps from input feature maps with Fourier-based low-frequency filtering, and subsequently masks edge-irrelated tokens with low energy values.
Extracting edge representations. Given the input feature maps $z_{i}\in\mathbb{R}^{N\times C}$, where $N$ is the number of tokens and $C$ is the number of channels. We first reshape the input features into $z_{i}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ are equal to $\lfloor\sqrt{N}\rfloor$. Then, we obtain its Fourier transformations:
@@ III-B Energy-based Edge Tokens Selection
The frequency signal $\mathcal{F}(z_{i})$ is then decomposed into the amplitude component $\mathcal{A}(z_{i})\in\mathbb{R}^{H\times W\times C}$ and the phase component $\mathcal{P}(z_{i})\in\mathbb{R}^{H\times W\times C}$, where the amplitude spectrum contains low-level statistics (e.g., style information) while the phase spectrum encodes high-level semantics (e.g., edge and shape). With the low-frequency components being shifted to the center of the amplitude spectrum, we employ a binary mask $\mathcal{M}_{r}\in\mathbb{B}^{H\times W}$, whose value is zero only for the center region of scale $r$, to controls the scale of low-frequency component to be discarded. Afterwards, we can obtain the high-frequency amplitude spectrum $\hat{\mathcal{A}}(z_{i})$:
@@
Selecting edge tokens by energy. Since $\hat{z}_{i}$ primarily retains edge-related information, to quantify the extent of edge information contained within each token, we calculate its activation value within $\hat{z}_{i}$, which is formulated as:
$E(\hat{z}_{i})_{\text{token}_{j}}=\frac{1}{C}\sum_{c=0}^{C-1}\hat{z}_{i}(j,c).$ (4)
The larger the activation value, the more edge information the token contains. With the token-level activation $E(\hat{z}_{i})_{\text{token}_{j}}$, we compute the $p$-th percentile, denoted as $q_{p}$, and perform a thresholding operation to obtain the edge-information guided mask $\mathcal{M}_{\text{edge}}\in\mathbb{B}^{N\times C}$. The element is set to $1$ if the corresponding element in $E(\hat{z_{i}})$ falls within the top $p$ percentage elements of $E(\hat{z}_{i})$, and set to 00 otherwise: $\mathcal{M}_{\text{edge}}(j,c)=\mathbb{I}(E(\hat{z}_{i})_{\text{token}_{j}}\geq q_{p}).$ Finally, we could distinguish and extract the edge tokens from the input features $z_{i}$ with the binary mask $\mathcal{M}_{\text{edge}}$, which is formulated as:
@@
Noting that in the process, the low-frequency filtered representation $\hat{z}_{i}$ is leveraged to generate the token-level mask $\mathcal{M}_{\text{edge}}$, which is applied to the original representation $z_{i}$ for preserving global shape features. The motivation behind this operation is that the low-frequency filtering could also lead to the loss of crucial semantic information surrounding edges, which may impede model training. Thanks to our proposed energy-based edge tokens selection (ETS), the augmented representations could retain critical object information within the edge-related tokens while discarding the edge-irrelevant tokens. We also evaluate the strategy that directly utilizes low-frequency filtering representation $\hat{z}_{i}$ to provide shape features in Sec. IV-E. The results validate the effectiveness of our ETS in preserving useful semantic information around object edges while suppressing model learning of domain-specific features.

### v2

# SETA: Semantic-Aware Edge-Guided Token Augmentation for Domain Generalization

@@ Abstract
Domain generalization (DG) aims to enhance the model robustness against domain shifts without accessing target domains. A prevalent category of methods for DG is data augmentation, which focuses on generating virtual samples to simulate domain shifts. However, existing augmentation techniques in DG are mainly tailored for convolutional neural networks (CNNs), with limited exploration in token-based architectures, i.e., vision transformer (ViT) and multi-layer perceptrons (MLP) models. In this paper, we study the impact of prior CNN-based augmentation methods on token-based models, revealing their performance is suboptimal due to the lack of incentivizing the model to learn holistic shape information. To tackle the issue, we propose the Semantic-aware Edge-guided Token Augmentation (SETA) method. SETA transforms token features by perturbing local edge cues while preserving global shape features, thereby enhancing the model learning of shape information. To further enhance the generalization ability of the model, we introduce two stylized variants of our method combined with two state-of-the-art (SOTA) style augmentation methods in DG. We provide a theoretical insight into our method, demonstrating its effectiveness in reducing the generalization risk bound. Comprehensive experiments on five benchmarks prove that our method achieves SOTA performances across various ViT and MLP architectures. Our code is available at https://github.com/lingeringlight/SETA.
@@ I Introduction
Among recent DG methods, data augmentation methods have exhibited promising performance [19, 11, 13, 14]. These methods primarily assume that domain gaps are caused by the disparity in style, thus introducing channel-level style perturbations for simulating potential domain shifts. While these methods have achieved notable success on CNN architectures, limited research investigates their effects on ViT or MLP models. ViT and MLP models have shown excellent generalization ability [13, 20], attributed to their capacity to learn spatial dependencies among tokens, which encode the structural shape information of objects, thus exhibiting a superior shape bias than CNNs [13, 15]. Current DG augmentation methods focus on perturbing style information, which has proven particularly effective for CNNs due to their severe texture bias. However, for token-based models (ViTs and MLPs) with inherently lower texture bias, it remains uncertain whether augmenting style information isolatedly could sufficiently enhance their generalization. Besides, previous methods mainly generate perturbations at the channel level, whereas token-based models focus on learning dependencies among tokens at the spatial level. Due to the lack of token-level augmentation for DG, the model could inevitably learn and extract domain-specific information from class-irrelevant tokens. Therefore, how to design an effective token-level augmentation for DG to enhance the generalization of token-based models sufficiently remains an open question.
@@
In this paper, we initially analyze the impact of existing SOTA augmentation methods in DG on the shape bias of the models. Specifically, we evaluate the robustness of these methods to diverse biased perturbations, focusing on two representative advanced methods: DSU [14], which is based on statistical perturbation, and ALOFT [13], which exploits frequency space augmentation. The experiments are conducted on the representative MLP-like model GFNet [18]. As shown in Fig. 1 (a), both DSU and ALOFT methods exhibit the ability to enhance the model generalization. The enhancement is attributed to their perturbations on style features, implicitly boosting the ability of the model to capture edge information, as indicated in Fig. 1 (b). Moreover, the enhanced edge-capturing ability also contributes to the increased model robustness against style perturbations, as depicted in Fig. 1 (c). However, as identified in [21, 22], image edges encompass not only global object shapes (i.e., category-related semantics), but also local spurious edges (i.e., category-irrelated edges like backbone textures). Since existing style augmentation methods for DG lack spatial token perturbation, they fail to suppress the domain-specific local edges, resulting in the limited ability of the model to perceive the holistic shape structures of objects, as proved in Fig. 1 (d). Consequently, the model inevitably overfits the tokens encoding domain-specific features, which severely impedes its generalizability to unseen domains. These observations prompt a critical question: Could we augment edge features at the token level to explicitly enhance the shape bias of the model, thus further improving the generalization ability of ViT and MLP architectures?
@@
To this end, we propose a novel Semantic-aware Edge-guided Token Augmentation (SETA) for DG, aiming at enhancing the generalization performance of ViT and MLP models by introducing token-level local edge perturbations. Unlike previous DG augmentation methods that focus on diversifying channel-level styles, our method concentrates on generating token-level edge perturbations to guide the model in emphasizing global shape features during training. Specifically, we initially design the Activation-based Edge Tokens Selection (ETS) module, which extracts edge maps from input features using low-pass filters and calculates the activation associated with each token to distinguish the edge-relevant tokens. Subsequently, random pairs of samples are selected from the batch, with one sample retaining only the edge tokens and the other being modified by the Shape Tokens Shuffling (STS) module that shuffles token locations within the sample. Finally, we conduct the random mixing of the two samples using either Mixup or CutMix, with the augmented sample being assigned the label of the sample providing edge tokens. Furthermore, we extend our method by integrating SOTA style augmentation techniques to enhance both shape sensitivity and robustness to style shifts. We evaluate our methods on five challenging datasets, employing various ViT and MLP architectures. The results prove the effectiveness of our method to enhance model generalization. Our contributions are summarized as follows:
@@
  • We theoretically prove that incentivizing shape sensitivity through our SETA could tighten the generalization risk bound and improve the generalizability of the model.
@@ II Related Work
Domain generalization. Domain generalization (DG) aims to learn generalized representations from source domains that generalize well to arbitrary unseen target domains. Existing DG methods address domain shift from various perspectives, including domain alignment [23, 24], meta-learning [9], data augmentation [25], ensemble learning [26, 27] and regularization methods [28, 29, 30]. The majority of relevant works to our method address the DG issue by data augmentation. These methods operate under the assumption that domain gaps predominantly arise from style shifts, thereby endeavoring to simulate domain shifts by generating diverse styles while keeping semantics consistency. Specifically, one prevalent line is to perturb the feature statistics (i.e., mean and variance) by sample interpolation [19, 31], modeling distribution [14] or learning methods [32, 33]. Moreover, some recent works propose to leverage the semantic-preserving property of the Fourier phase component and introduce diverse noise into the amplitude spectrum to generate stylized data [34, 35]. However, these methods are mainly designed for CNN architectures, which focus on perturbing channel-level style features. When applied to ViT or MLP models that learn spatial dependencies among different tokens, as these methods lack token-level augmentation, the model inevitably overfits domain-specific tokens that encode spurious local edge information, which impedes its generalizability. Although some pioneering works [36, 37] have recently explored ViT-based methods for DG, these methods rely on supervisory signals from additional tokens to implicitly constrain the model learning of domain-invariant features, ignoring the presence of domain-related information within feature tokens. Due to the lack of explicit constraints on feature tokens, when facing significant domain shifts, these tokens are likely to contain substantial domain-specific information, leading to overfitting to source domains. Therefore, we propose a token-level augmentation method that explicitly enhances domain-invariant feature extraction within feature tokens, which is the first DG method to improve the generalization ability of the model from the perspective of incentivizing global shape bias.
@@
Figure 2: An overview of the proposed SETA. Our SETA consists of three core modules, including 1) Activation-based Edge Tokens Selection (ETS) that distinguishes and extracts tokens containing edge information; 2) Shape Tokens Shuffling (STS) that generates texture noise by shuffling tokens from another sample, which disrupts holistic shape while keeping local edges; 3) Token Mixing module that superposes edge tokens from an object sample onto the shuffled tokens from another sample. The augmented sample is assigned the label of the sample contributing the edge tokens. We design two stylized variants of SETA, utilizing SOTA DG augmentation methods, i.e., DSU [14] and ALOFT [13], to create stylized token-shuffled samples to simulate potential domain shifts.
@@
Data augmentation for improving shape bias. Previous studies have revealed that compared to human vision, CNNs are revealed to be more sensitive to image texture for object classification, indicating that the trained models lack a full understanding of different object categories [22, 38]. To reduce this disparity, augmentation methods aim to perturb texture representations, which could explicitly reduce the texture bias and implicitly boost the shape bias of the model [22]. Several works also indicate that an increased shape bias in CNNs is associated with improved model robustness [39, 40]. The exploration of texture and shape biases has extended to ViTs and MLPs, showing that they exhibit a lower texture bias compared to CNNs [15]. It is attributed to the ability of token-based models to effectively learn dependencies among tokens, thereby enhancing their capability to capture global semantic information. Some advanced token-level augmentations have been explored to improve the learning capacity of ViTs. TokenMixup [41] utilizes attention maps within ViTs to construct a saliency-aware token mixup for supervised classification, aiming to help the model learn label-related class information from tokens. PMTrans [42] employs a game-theory-based learnable beta distribution to sample and mix patches from source and target domains, aligning the labeled source domain with the unlabeled target domain. However, in DG scenarios, these methods cannot guarantee the model’s performance on unseen target domains. Due to the lack of explicit constraints on domain-specific texture information, these methods inevitably suffer from texture bias, significantly impairing the model’s ability to generalize across different domains. Although some DG methods [19, 34] attempt to perturb style information while keeping the others unchanged (e.g., edges), they fail to suppress spurious edge information (i.e., shape-irrelevant textures), limiting their ability to enhance the shape bias. In the paper, we propose a semantic-aware edge-guided token augmentation method, which can discern and perturb spurious edges while retaining crucial shape information. Our method compels the model to learn a comprehensive understanding of object shape from the edge-perturbed representations, thereby enhancing its shape bias effectively.
@@ III-A Setting and Overview
In standard DG model training, the model is directly trained to predict the label of the original samples. However, as emphasized in [38, 40, 43], this training method could induce sensitivity to style and texture in source domains, consequently hindering model generalization on unseen target domains. To tackle this issue, we propose a novel token-level augmentation method, namely Semantic-aware Edge-guided Token Augmentation (SETA), for ViT and MLP models. The key idea of SETA is to generate shape-enhanced representations by mixing edge-related tokens from a sample and shape-disrupted tokens from other samples. This operation compels the model to distinguish the overall shape of objects from the shape-irrelated edge-perturbed representations, thereby enhancing its sensitivity to shape information. To further enhance the model robustness to domain shifts, we explore a paradigm that integrates SETA with SOTA augmentation methods in DG, yielding two stylized variants of SETA. An overview of our method is depicted in Fig. 2. Below we introduce each component of our method and provide a theoretical insight into its effectiveness for improving model generalization.
III-B Activation-based Edge Tokens Selection
Previous works [21, 22] have revealed that image edges encompass both global shape structures (i.e., category-related object shapes) and local edge textures (i.e., category-irrelevant local edges). However, directly distinguishing shape concepts from images is challenging. To address this, our method focuses on perturbing local edge textures to implicitly enhance the model’s ability to learn the remaining shape structure information. Specifically, following [13, 35], we first design the Activation-based Edge Token Selection (ETS) module to extract edge information, i.e., leveraging Fourier transformation to extract high-frequency component of representation.
Extracting edge representations. Given the input feature map $z_{i}\in\mathbb{R}^{N\times C}$, where $N$ is the number of tokens and $C$ is the number of channels. We first reshape the input feature into $z_{i}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ are equal to $\lfloor\sqrt{N}\rfloor$. Then, we obtain its Fourier transformation:
@@ III-B Activation-based Edge Tokens Selection
The frequency signal $\mathcal{F}(z_{i})$ is then decomposed into the amplitude component $\mathcal{A}(z_{i})\in\mathbb{R}^{H\times W\times C}$ and the phase component $\mathcal{P}(z_{i})\in\mathbb{R}^{H\times W\times C}$, where the amplitude spectrum contains low-level statistics (e.g., style information) while the phase spectrum encodes high-level semantics (e.g., edge and shape). With the low-frequency components being shifted to the center of the amplitude spectrum, we employ a binary mask $\mathcal{M}_{r}\in\mathbb{B}^{H\times W}$, whose value is zero only for the center region of scale $r$, to control the scale of low-frequency component to be discarded. Afterwards, we can obtain the high-frequency amplitude spectrum $\hat{\mathcal{A}}(z_{i})$:
@@
Selecting edge tokens by activation. Since $\hat{z}_{i}$ primarily retains edge-related information, to quantify the extent of edge information contained within each token, we calculate its activation value within $\hat{z}_{i}$, which is formulated as:
$A(\hat{z}_{i})_{\text{token}_{j}}=\frac{1}{C}\sum_{c=0}^{C-1}\hat{z}_{i}(j,c).$ (4)
The larger the activation value, the more edge information the token contains. With the token-level activation $A(\hat{z}_{i})_{\text{token}_{j}}$, we compute the $p$-th percentile, denoted as $q_{p}$, and perform a thresholding operation to obtain the edge-information guided mask $\mathcal{M}_{\text{edge}}\in\mathbb{B}^{N\times C}$. The element is set to $1$ if the corresponding element in $A(\hat{z_{i}})$ falls within the top $p$ percentage elements of $A(\hat{z}_{i})$, and set to 00 otherwise: $\mathcal{M}_{\text{edge}}(j,c)=\mathbb{I}(A(\hat{z}_{i})_{\text{token}_{j}}\geq q_{p}).$ Finally, we could distinguish and extract the edge tokens from the input features $z_{i}$ with the binary mask $\mathcal{M}_{\text{edge}}$, which is formulated as:
@@
Noting that in the process, the low-frequency filtered representation $\hat{z}_{i}$ is leveraged to generate the token-level mask $\mathcal{M}_{\text{edge}}$, which is applied to the original representation $z_{i}$ for preserving global shape features. The motivation behind this operation is that the low-frequency filtering could also lead to the loss of crucial semantic information surrounding edges, which may impede model training. Thanks to our proposed activation-based edge tokens selection (ETS), the augmented representations could retain critical object information within the edge-related tokens while discarding the edge-irrelevant tokens. We also evaluate the strategy that directly utilizes low-frequency filtering representation $\hat{z}_{i}$ to provide shape features in Sec. V. The results validate the effectiveness of our ETS in preserving useful semantic information around object edges while suppressing model learning of domain-specific features.